If 3 is added to a certain number and the sum multiplied by 2, the result is the same as nine subtracted from three times the number. Find this number.
Answer by Earlsdon(6103)

Let n be the number, then from the problem description, you can write:
"Two times the sum of the number and three is the same as (=) nine subtracted from three times the number."
2(n+3) = 3n-9 Simplify.
2n+6 = 3n-9 Subtract 2n from both sides.
6 = n-9 Add 9 to both sides.
15 = n
The number is 15.